# Sat2Flow: A Structure-Aware Diffusion Framework for Human Flow Generation from Satellite Imagery

### v1

@@ Abstract
Origin-Destination (OD) flow matrices are essential for urban mobility analysis, underpinning applications in traffic forecasting, infrastructure planning, and policy design. However, existing methods suffer from two critical limitations: (1) reliance on auxiliary features (e.g., Points of Interest, socioeconomic statistics) that are costly to collect and have limited spatial coverage; and (2) sensitivity to spatial topology, where minor index reordering of urban regions (e.g., census tract relabeling) disrupts structural coherence in generated flows. To address these challenges, we propose Sat2Flow, a latent structure-aware diffusion-based framework that generates structurally coherent OD flows using solely satellite imagery as input. Our approach introduces a multi-kernel encoder to capture diverse regional interactions and employs a permutation-aware diffusion process that aligns latent representations across different regional orderings. Through a joint contrastive training objective that bridges satellite-derived features with OD patterns, combined with equivariant diffusion training that enforces structural consistency, Sat2Flow ensures topological robustness under arbitrary regional reindexing. Experimental results on real-world urban datasets demonstrate that Sat2Flow outperforms both physics-based and data-driven baselines in numerical accuracy while preserving empirical distributions and spatial structures under index permutations. Sat2Flow offers a globally scalable solution for OD flow generation in data-scarce urban environments, eliminating region-specific auxiliary data dependencies while maintaining structural invariance for robust mobility modeling.
Codehttps://github.com/5hawnWang/Sat2Flow
@@ Introduction
The proposed Sat2Flow framework is implemented in PyTorch using the AdamW optimizer with a learning rate of $10^{-3}$ and weight decay of $10^{-2}$. Contrastive learning and diffusion learning are trained for 1000 epochs with a batch size of 12. For data augmentation, we apply random permutations to regional indices at 50% intensity to enhance robustness to index reordering. Satellite imagery is encoded into 768-dimensional features, while the multi-kernel encoder is configured with kernel dimension 64 and 8 kernels. All experiments were performed on a single NVIDIA GeForce RTX A6000 GPU with 50GB of memory.
@@ Acknowledgments
This research was supported in part by Shenzhen Science and Technology Program (No. JCYJ20240813113300001, 20231127180406001, JYCJ20220530152817039, JCYJ20240813113218025).

### v2

@@ Abstract
Origin-Destination (OD) flow matrices are critical for urban mobility analysis, supporting traffic forecasting, infrastructure planning, and policy design. Existing methods face two key limitations: (1) reliance on costly auxiliary features (e.g., Points of Interest, socioeconomic statistics) with limited spatial coverage, and (2) fragility to spatial topology changes, where reordering urban regions disrupts the structural coherence of generated flows. We propose Sat2Flow, a structure-aware diffusion framework that generates structurally coherent OD flows using only satellite imagery. Our approach employs a multi-kernel encoder to capture diverse regional interactions and a permutation-aware diffusion process that maintains consistency across regional orderings. Through joint contrastive training linking satellite features with OD patterns and equivariant diffusion training enforcing structural invariance, Sat2Flow ensures topological robustness under arbitrary regional reindexing. Experiments on real-world datasets show that Sat2Flow outperforms physics-based and data-driven baselines in accuracy while preserving flow distributions and spatial structures under index permutations. Sat2Flow offers a globally scalable solution for OD flow generation in data-scarce environments, eliminating region-specific auxiliary data dependencies while maintaining structural robustness for reliable mobility modeling.
Codehttps://github.com/ai4city-sztu/Sat2Flow
@@ Introduction
The proposed Sat2Flow framework is implemented in PyTorch using the AdamW optimizer with a learning rate of $10^{-3}$ and weight decay of $10^{-2}$. Contrastive learning and diffusion learning are trained for 1000 epochs with a batch size of 12. For data augmentation, we apply random permutations to regional indices at 50% intensity to enhance robustness to index reordering. Satellite imagery is encoded into 768-dimensional features, while the MK encoder is configured with kernel dimension 64 and 8 kernels. All experiments were performed on a single NVIDIA GeForce RTX A6000 GPU with 48GB of memory.
@@ Acknowledgments
This research was supported by Shenzhen Science and Technology Program (No. JCYJ20240813113300001, 20231127180406001).